40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 7 23:24:41 2014, Jenne, Update, General, IFO plan, list o' things to do 
    Reply  Wed Jan 8 13:46:39 2014, Jenne, Update, General, IFO plan, PRM baffle PRMcage_6Sept2012.JPG
    Reply  Wed Jan 8 17:53:26 2014, manasa, Update, General, IFO plan, IPANG telescope IPANG_layout.pngIPANG.png
       Reply  Thu Jan 9 17:21:45 2014, rana, Update, General, IFO plan, IPANG telescope 
    Reply  Thu Jan 9 10:34:58 2014, Steve, Update, PSL, PSL pointing changes in pitch Lfnfdnc.png24hPSLpointing.png
       Reply  Fri Jan 10 10:28:03 2014, Steve, Update, PSL, PSL pointing changes  clamped.png
Message ID: 9542     Entry time: Thu Jan 9 10:34:58 2014     In reply to: 9534     Reply to this: 9545
Author: Steve 
Type: Update 
Category: PSL 
Subject: PSL pointing changes in pitch 

  IOO QPDs tested in dark, lighted and open PSL enclosure. The created temperature change 0.03 C has  effect on monitoring  in pitch.

 

 Atm1,  all lights off 10 min, PSL enclosure lights on  10 min, all lights off 15 min, open  door # 11 at north east corner of enclosure ( HEPA filters are running at 30V ) for 10 min, closed-dark enclosure 15 min

              dark 10, lighted 10, dark 15, open-dark 10 and closed-dark 15 minutes

 

Atm2, Pitch drift of 24 hours does not recover

Attachment 1: Lfnfdnc.png  44 kB  | Hide | Hide all
Lfnfdnc.png
Attachment 2: 24hPSLpointing.png  54 kB  | Hide | Hide all
24hPSLpointing.png
ELOG V3.1.3-